A discharging circuit has a voltage across the capacitor of v = V e^((-t)/RC) Determine dv/dt, the rate of change in voltage in the discharging circuit capacitor. Then find dv/dt when t = T.
1
Expert's answer
2020-04-29T17:43:46-0400

Given "v = V e^{-t\/RC}" .

So, "\frac{dv}{dt} = V \frac{-1}{RC} e^{-t/RC} =-\frac{V}{RC} e^{-t/RC}" .


Now "dv\/dt" at t=T is "\frac{dv}{dt}|_{t=T} =-\frac{V}{RC} e^{-T/RC}"